Each unit of grain A contains 2 grams of fat and 80 calories, and unit B has 3 grams of fat and 60 calories. If the final product is to provide 18 grams of fat and 480 calories, how much of each type of grain should be used. I think A is 3 and B is 4, but I don't know how to show it.
Answer by mananth(16946)

Unit A------------- x Numbers
Unit B ------------ y Numbers
2x+3y=18 .............1
80x+60y=480.............2
Eliminate y
multiply (1)by -20
Multiply (2) by 1
-40x-60 y=-360
80x+60y=480
Add the two equations
40x=120
/40
x=3 Unit A
plug value of x in (1)
2x+3y=18
6+3y= 18
3y=18-6
3y=12
y=4 Unit B